OpenStudy (anonymous):

A town's population has been growing linearly. In the beginning of 2000, the population was 36000, and the population in 2006 is 49200. Write an equation for the population of the town where x represents the number of years after the beginning of 2000. P(x)= What is/was/will be the population in the 2009? How many years from the beginning of 2000 did/will it take the population to reach 75800?

OpenStudy (campbell_st):

well you need to find the slope of the line change in population/change in time the 2 points are (2000, 36000) and (2006, 49200) when you get the slope, m then you will have an equation y = mx + b use either point to find b

OpenStudy (anonymous):

ah okay thanks, so y2-y1/x2-x1 to find the slope?

OpenStudy (campbell_st):

yep thats it... populations on top... time in the denominator

OpenStudy (anonymous):

i get 12800/6 then plug it into an equation and get y=12800/6x+b to find be i solve for it therefore getting 36000=12800/6(0)+b 0 because x represents the number of years

OpenStudy (anonymous):

therefore b=36000

OpenStudy (campbell_st):

seems to make sense.... looking at the calculations.

OpenStudy (anonymous):

Inputted the calculations and got it wrong

OpenStudy (anonymous):

p/x=12800/6x+36000

OpenStudy (campbell_st):

well isn't the change in population 49200 - 36000 =13200 which is divisible by 6 so m = 2200 so the equation is y = 2200x + b so b is still 36000 hence the equation is y = 2200x + 36000

OpenStudy (anonymous):

ohh I did y2-y1/x2-x1 So I didn't need to do that then?

OpenStudy (campbell_st):

well its the same thing..... you just had the change in population as 12800 when it should be 13200. the rest is correct... I just read you solution and thought... 12800/6 seemed odd... in that is wasn't an integer answer

OpenStudy (anonymous):

oh okay, makes sense math was off then, so now to figure out for 75,800 population I plug in 75800 in y and solve for x?

OpenStudy (campbell_st):

yes and for the 2009... plug in x = 9

OpenStudy (anonymous):

i got the 2009 one but the 75800 i get a decimal, i get 18.0909090....